You planted a 16 inch hemlock tree that grows at a rate of 4 inches per year. Then 5 years later you planted a 10inch blue spruce that grows at a rate of 6 inches per year. In how many years after you planted the hemlock tree will the two trees be the same height? How tall will the trees be?
Answer by williamson.family05(2)

If the 16 inch hemlock tree grows 4 inches per year then in 1 year the tree would be 20 inches tall, the second year it would be 24 inches tall, and the third year it would be 28 inches tall.
1 year = 16+4 = 20
2 year = 20+4 = 24
3 year = 24+4 = 28
If the 10 inch spruce grows 6 inches per year then in 1 year the tree would be 16 inches tall, the second year it would be 22 inches tall, and the third year it would be 28 inches tall.
1 year = 10+6 = 16
2 year = 16+6 = 22
3 year = 22+6 = 28
So in 3 years the trees will be the same height. Height = 28 inches